OpenStudy (anonymous):

how to calculate length of column given 2 peak with retention time 92 and 107s, using a column with an H of1.6cm/plate.

OpenStudy (aaronq):

they didn't give you the width of the peak?

OpenStudy (aaronq):

the only way i can think of this is \(H=\dfrac{L}{N}=\dfrac{L}{5.55(\dfrac{99.5~s}{w})^2}\)

OpenStudy (anonymous):

there was no width of the peak

OpenStudy (anonymous):

1.6=L(what im looking for) / N (which is not given) to find N ill use you equation 5.55... but there is width. i read though the whole chapter could not find anything that can omit the width

OpenStudy (aaronq):

yeah, i don't know any other way :S

OpenStudy (anonymous):

exact question: What length of column would be required to just resolve 2 peak with retention times of 92 and 107s, respectively, using a column with an H of 1.60cm/plate.